what is the solution of the following system of linear equations?
x + 2y = -7
x = y + 23
a. (11,-10)
b. (13, -10)
c. (13, 10)
d. (10, -13)

Answer by mananth(16946)

x + 2y = -7
x = y + 23
..
x+2y=-7
x-y=23
subtract one from another to eliminate x
x+2y-x+y=-7-23
3y=-30
y=-10
..
Plug the value of y in any equation above
x-(-10)=23
x+10=23
x=13